A store has clearance items that have been marked down by 45%. They are having a sale, advertising an additional 55% off clearance items. What percent of the original price do you end up paying?

Answer by ikleyn(52786)

.

After first discount, the sale price per item is  1-0.45 = 0.55 of its original price.



After second discount, the customer pays  (1-0.55)*0.55 = 0.45*0.55 = 0.2475 of the original price, 

or  24.75% of the original price.    ANSWER

Solved.